# Thermal quasi-particle theory

### v1

@@ VI Thermal quasi-particle energies
$10^{5}$ $-0.07423$ $-0.05604$ $0.04959$
$10^{6}$ $-0.56092$ $-0.56475$ $-0.44097$
$10^{7}$ $-3.26523$ $-3.26554$ $-3.17327$
$10^{8}$ $-4.92771$ $-4.92775$ $-4.91206$
Table 7 compares the slope $\partial U/\partial N$ at $N=\bar{N}$ computed analytically by Eq. (52) for thermal HF and QP(2) theories against the slope of thermal FCI theory obtained by numerical differentiation. The three sets of values are in reasonable agreement with one another at all $T$ considered. At low $T$, thermal QP(2) theory is in much better agreement with thermal FCI theory than thermal HF theory as the former takes into account electron correlation. At higher $T$, thermal QP(2) and HF theories are more alike than thermal FCI theory, but they all seem to converge at the same high $T$ limit.
To summarize the results of the foregoing analysis, the orbital energies $\epsilon_{p}$ of thermal quasi-particle theories (including thermal HF theory) cannot be directly related to the ionization or electron-attachment energy from/into the $p$th orbital except at $T=0$. This is simply because an ionization in an ideal gas of molecules at $T\gg 0$ is not a removal of an electron from the $p$th orbital in every molecule; rather, it is a removal of an electron from various orbitals of all molecules at some probabilities (even in a strict independent-particle picture). This is, in turn, equivalent to a removal of some fraction of an electron from various orbitals of a single average molecule. The $\epsilon_{p}$ at $T>0$ signifies the increase in internal energy upon adding an infinitesimal fraction of an electron in the $p$th spinorbital of this “average” molecule. This quantity may not directly correspond to any observable, but it can be combined to produce a macroscopic thermodynamic observable, for which thermal QP(2) theory is in better agreement with thermal FCI theory than thermal HF theory.
VII Conclusions
The thermal quasi-particle theory has been introduced. It is a quasi-independent-particle theory and a finite-temperature extension of the widely used thermal HF theory. Its entropy and chemical-potential formulas are of the one-electron type,
@@ VII Conclusions
the theory satisfies all thermodynamic relations and the $\Omega$ is a variational minimum with respect to $f_{p}^{-}$.
Fermi–Dirac and thermal HF theories can be viewed as the zeroth- and first-order instances, respectively, in the hierarchy of thermal quasi-particle theories. It can be viewed as a diagonal, frequency-independent approximation to a finite-temperature extension of the inverse Dyson equation, which is an exact one-particle theory.Hirata et al. (2017) In other words, the $\epsilon_{p}$ thus defined is a finite-temperature extension of the Dyson self-energy in the diagonal, frequency-independent approximation. Generally, the thermal self-energy is shown to suffer from the severe kind of Kohn–Luttinger nonconvergence problem and so it is inappropriate to include anomalous-diagram terms in the definition of approximate $U$ or $\Omega$.
Our preliminary implementation and comparison with the thermal FCI benchmarks suggests that thermal QP(2) theory performs distinctly better than thermal HF theory at low $T$ and may even outperform finite-temperature MBPT(2) at intermediate $T$ by including electron-correlation effects in the orbital energies. At higher $T$, the entropy contribution dominates and the differences among various methods become relatively insignificant.
The direct physical meaning of $\epsilon_{p}$ is the increase in internal energy upon infusion of an infinitesimal fraction of an electron in the $p$th spinorbital. It does not correspond to the ionization energy or electron-attachment energy of a molecule at $T>0$. However, it can be used as ingredients from which electron-correlated thermodynamic observables can be computed.
@@ Appendix A Kohn–Luttinger nonconvergence
where subscript “$L$” means that only diagrammatically linked contributions are to be retained, and the last term (carrying a $\beta$ multiplier) is the anomalous-diagram term. This is still far from the whole second-order internal energy [Eq. (74) of Ref. Hirata and Jha, 2020], but the presence of the last anomalous-diagram term will make the following analysis meaningful.
@@
where the right-hand side is Dirac’s $\delta$ function, which is divergent when the chemical potential $\mu$ coincides with one of the orbital energies.Kohn and Luttinger (1960) Therefore, a Kohn–Luttinger nonconvergence occurs when $F_{pp}\neq 0$ and at the same time, the reference wave function is degenerate (i.e., the HOMO and LUMO have the same energy as $\mu$). While the divergence of this term can still be avoided by adopting a HF reference wherein $F_{pq}=0$,Luttinger and Ward (1960) the fourth term remains divergent for $p=r$ and $q=s$ in a degenerate reference.Hirata (2021b, 2022) It should be recalled that the second-order correction to the ground-state energy is always finite regardless of whether the zeroth-order ground state is degenerateHirschfelder and Certain (1974) or nondegenerate.Shavitt and Bartlett (2009) Hence, $U^{(2)}$ is not convergent at the correct zero-temperature limit, which is always finite.
Figure 11: Second-order thermal Dyson self-energy diagrams of Eq. (61), which include anomalous diagrams with no resolvent lines. They are obtained by cutting an edge in each of the internal energy diagrams of Fig. 10 and trimming the dangling lines.
As per Eq. (44), the corresponding thermal self-energy is obtained by taking the derivative of $U^{(2)}$ with respect to $f_{p}^{-}$, i.e.,
@@
$=$ $\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}-\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{q}^{(0)}-\epsilon_{p}^{(0)}}f_{q}^{-}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{\langle qr||pr\rangle F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{F_{qp}\langle pr||qr\rangle}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{p}^{-}f_{q}^{+}$
$-{\beta}\sum_{q}^{\text{denom.}=0}F_{qp}F_{pq}f_{q}^{+}+{\beta}\sum_{q}^{\text{denom.}=0}F_{qp}F_{pq}f_{q}^{-}$
$-{\beta}\sum_{q,r}^{\text{denom.}=0}{\langle qr||pr\rangle F_{pq}f_{p}^{-}f_{q}^{+}}$
$-{\beta}\sum_{q,r}^{\text{denom.}=0}{F_{qp}\langle pr||qr\rangle f_{p}^{-}f_{q}^{+}}$
@@
The corresponding diagrams are shown in Fig. 11, which are obtained by opening each of the $U^{(2)}$ diagrams in Fig. 10 by deleting an edge.
This formula suffers from a more pervasive kind of the Kohn–Luttinger nonconvergence problem. Take the penultimate term for example. It sums over the $p=r$ and $q=s$ case, whose summand diverges as $T\to 0$ ($\beta\to\infty$) for any reference—degenerate or nondegenerate—because a $f_{p}^{-}$ factor is no longer there to create a $\delta$ function [Eq. (60)] which diverges only when $\mu=\epsilon_{p}^{(0)}$. In other words, a finite-temperature generalization of the self-energy, when it is derived more rigorously from finite-temperature MBPT and thus includes anomalous-diagram terms, is far more prone to divergence as $T\to 0$ and will be uselessly erroneous at low $T$. This is not surprising because in thermodynamics all states with any number of electrons are summed over in the partition function and they always include numerous exactly degenerate zeroth-order states.
This finding may bode ill for the Luttinger–Ward functionalLuttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik et al. (2015); Rossi and Werner (2015); Welden et al. (2016); Gunnarsson et al. (2017); Lin and Lindsey (2018) because opening a diagram of this functional will result in a thermal self-energy or Green’s function diagram that may always suffer from the same severe low-temperature breakdown.
@@ Appendix B Feynman–Dyson nonconvergence
In Ref. Hirata et al., 2024 was reported that a Feynman–Dyson perturbation expansion of the Dyson self-energy is always divergent in many frequency domains. Therefore, even apart from the divergence of the thermal self-energy caused by anomalous-diagram terms analyzed in Appendix A, the zero-temperature self-energy is expected to be divergent or at least excessively erroneous for low- and high-lying states. This is alleviated (or hidden from view) by the frequency-independent approximation inherent in the ansatz of thermal quasi-particle theory, but in a solid, whose self-energy forms continuous energy bands, we can no longer expect to be able to avoid this deep-rooted pathology.Hirata et al. (2024)
We therefore expect that in a solid-state implementation of thermal QP(2) theory (which is underway in our laboratory), we will introduce active orbitals (or active energy bands) and define and evaluate portions of $\langle E^{(2)}\rangle$ and of $\Sigma_{pp}^{(2)}$ involving these orbitals. These active orbitals are the ones whose zeroth-order energies fall within the so-called “central overlapping bracket” where the Feynman–Dyson perturbation expansion is guaranteed to have a nonzero radius of convergence.Hirata et al. (2024) Comparative performance of this method for solids will be presented in the future.

### v2

@@ III.2 Thermal quasi-particle energies
$10^{5}$ $-0.07423$ $-0.04741$ $0.04959$
$10^{6}$ $-0.56092$ $-0.55587$ $-0.44097$
$10^{7}$ $-3.26523$ $-3.26425$ $-3.17327$
$10^{8}$ $-4.92771$ $-4.92757$ $-4.91206$
Table 7 compares the slope $\partial U/\partial N$ at $N=\bar{N}$ computed analytically by Eq. (60) for thermal HF and QP(2) theories against the slope of thermal FCI theory obtained by numerical differentiation. The three sets of values are in reasonable agreement with one another at all $T$ considered. At low $T$, thermal QP(2) theory is in much better agreement with thermal FCI theory than thermal HF theory. This is because the former takes into account electron correlation. At higher $T$, thermal QP(2) and HF theories are more alike than thermal FCI theory, but they all seem to converge at the same high $T$ limit.
To summarize the results of the foregoing analysis, the orbital energies $\epsilon_{p}^{\text{QP}}$ of thermal quasi-particle theories (including thermal HF theory) cannot be directly related to the ionization or electron-attachment energy from/into the $p$th orbital except at $T=0$. This is simply because an ionization in an ideal gas of molecules at $T\gg 0$ is not a removal of an electron from the $p$th orbital in every molecule; rather, it is a removal of an electron from various orbitals of all molecules at some probabilities (even in a strict independent-particle picture). This is, in turn, equivalent to a removal of some fraction of an electron from various orbitals of an “average” molecule. The $\epsilon_{p}^{\text{QP}}$ at $T>0$ signifies the increase in internal energy upon adding an infinitesimal fraction of an electron in the $p$th spinorbital of this average molecule. This interpretation may be called thermal Janak’s theorem.Janak (1978) This quantity may not directly correspond to any observable, but it is likely combined to approximate macroscopic thermodynamic observables, for which thermal QP(2) theory is expected to be more accurate than thermal HF theory. A further analysis is needed to justify the use of these quasi-particle energies in characterizing, e.g., metal-insulator transitions, however.
IV Conclusions
The thermal quasi-particle theory has been introduced. It is a quasi-independent-particle theory and an electron-correlated extension of the widely used thermal HF theory. It is size-consistent and thus directly applicable to electron-correlated energy bands and thermodynamic functions of semiconductors and insulators at finite temperature.
Its entropy and chemical-potential formulas are of the one-electron type,
@@ IV Conclusions
the theory satisfies all fundamental thermodynamic relations and the $\Omega$ is a variational minimum with respect to $f_{p}^{-}$.
Fermi–Dirac and thermal HF theories can be viewed as the zeroth- and first-order instances, respectively, of the thermal quasi-particle theory hierarchy. It can be viewed as a diagonal, frequency-independent approximation to a finite-temperature extension of the inverse Dyson equation, which at zero temperature is an exact one-particle theory.Hirata et al. (2017) In other words, the $\epsilon_{p}$ thus defined is a finite-temperature extension of the Dyson self-energy in the diagonal, frequency-independent approximation.
It has also been revealed that the thermal self-energy suffers from a severe Kohn–Luttinger nonconvergence problem, making it inappropriate to include anomalous-diagram terms in the definition of approximate $U$ or $\Omega$. This may mean that the thermal quasi-particle theory hierarchy is not convergent at exactness, unfortunately.
Our preliminary implementation and comparison with the thermal FCI benchmarks suggests that thermal QP(2) theory performs distinctly better than thermal HF theory at low $T$ and may even outperform finite-temperature MBPT(2) at intermediate $T$ by including electron-correlation effects in the orbital energies. At higher $T$, the entropy contribution dominates and the differences among various methods become relatively insignificant. The zero-temperature limit of thermal QP(2) theory is MBPT(2) for internal energy and MBGF(2) (in the diagonal, frequency-independent approximation) or $\Delta$MP2 for orbital energies.
The direct physical meaning of $\epsilon_{p}$ is the increase in the internal energy upon infusion of an infinitesimal fraction of an electron in the $p$th spinorbital, i.e., thermal Janak’s theorem.Janak (1978) It does not correspond to the ionization energy or electron-attachment energy of a gas of molecules or a solid at $T>0$. However, it can be used as the key ingredients from which electron-correlated thermodynamic observables can be computed.
@@ Appendix A Kohn–Luttinger nonconvergence
where subscript “$L$” means that only diagrammatically linked contributions are to be retained, and the last term (carrying a $\beta$ multiplier) is the anomalous-diagram term. This is still far from the whole second-order internal energy [Eq. (74) of Ref. Hirata and Jha, 2020], but the presence of the last anomalous-diagram term will make clear that this ansatz is unworkable.
@@
where the right-hand side is Dirac’s delta function, which is divergent when the chemical potential $\mu$ coincides with one of the orbital energies.Kohn and Luttinger (1960) Therefore, a Kohn–Luttinger nonconvergence occurs when $F_{pp}\neq 0$ and at the same time, the reference wave function is degenerate (i.e., the HOMO and LUMO have the same energy as $\mu$). While the divergence of this term can be avoided by adopting a HF reference wherein $F_{pq}=0$,Luttinger and Ward (1960) the fourth term remains divergent for $p=r$ and $q=s$ in a degenerate reference.Hirata (2021b, 2022) It should be recalled that the second-order correction to the ground-state energy is always finite regardless of whether the zeroth-order ground state is degenerateHirschfelder and Certain (1974) or nondegenerate.Shavitt and Bartlett (2009) Hence, $U^{(2)}$ is not convergent at the correct zero-temperature limit, which is always finite.
Figure 11: Second-order thermal Dyson self-energy diagrams of Eq. (69), which include anomalous diagrams with no resolvent lines. They are obtained by deleting an edge in each of the internal energy diagrams of Fig. 10.
As per Eq. (42), the corresponding thermal self-energy is obtained by taking the derivative of $U^{(2)}$ with respect to $f_{p}^{-}$, leading to
@@
$=$ $\sum_{q}^{\text{denom.}\neq 0}\frac{F_{qp}F_{pq}}{\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}}f_{q}^{+}-\sum_{q}^{\text{denom.}\neq 0}\frac{F_{pq}F_{qp}}{\epsilon_{q}^{(0)}-\epsilon_{p}^{(0)}}f_{q}^{-}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{\langle qp||rp\rangle F_{rq}}{\epsilon_{r}^{(0)}-\epsilon_{q}^{(0)}}f_{r}^{-}f_{q}^{+}$
$+\sum_{q,r}^{\text{denom.}\neq 0}\frac{F_{qr}\langle rp||qp\rangle}{\epsilon_{r}^{(0)}-\epsilon_{q}^{(0)}}f_{r}^{-}f_{q}^{+}$
$-{\beta}\sum_{q}^{\text{denom.}=0}F_{qp}F_{pq}f_{q}^{+}+{\beta}\sum_{q}^{\text{denom.}=0}F_{pq}F_{qp}f_{q}^{-}$
$-{\beta}\sum_{q,r}^{\text{denom.}=0}{\langle qp||rp\rangle F_{rq}f_{r}^{-}f_{q}^{+}}$
$-{\beta}\sum_{q,r}^{\text{denom.}=0}{F_{qr}\langle rp||qp\rangle f_{r}^{-}f_{q}^{+}}$
@@
The corresponding diagrams are shown in Fig. 11, which are obtained by opening each of the $U^{(2)}$ diagrams in Fig. 10 by deleting an edge. Note the presence of numerous anomalous diagrams.
This formula suffers from a more pervasive kind of the Kohn–Luttinger nonconvergence problem. Take the last term for example. It sums over a set of indexes whose fictitious denominator $\epsilon_{r}^{(0)}+\epsilon_{s}^{(0)}-\epsilon_{p}^{(0)}-\epsilon_{q}^{(0)}$ is zero. It includes the $p=r$ and $q=s$ case, whose summand diverges as $T\to 0$ ($\beta\to\infty$) for any reference—degenerate or nondegenerate—because a $f_{p}^{+}$ factor is no longer there to create a delta function [Eq. (68)] which diverges only when $\mu=\epsilon_{p}^{(0)}$. In other words, a finite-temperature generalization of the self-energy containing anomalous-diagram terms is always divergent as $T\to 0$ and will be uselessly erroneous at low $T$.
This is not surprising because in thermodynamics all states with any number of electrons are summed over in the partition function and they include numerous exactly degenerate zeroth-order states. This is why we purposefully neglected anomalous-diagram terms when defining $U^{\text{QP(2)}}$. Generally, it is inappropriate to include them in the internal energy or grand potential, when their (closed) diagrams will be opened to form some kind of potentials.
This bodes ill for the Luttinger–Ward functionalLuttinger and Ward (1960); Baym and Kadanoff (1961); Baym (1962); Kozik, Ferrero, and Georges (2015); Rossi and Werner (2015); Welden, Rusakov, and Zgid (2016); Gunnarsson et al. (2017); Lin and Lindsey (2018) because opening a diagram of this functional may also result in thermal self-energy and Green’s function diagrams that suffer from the same severe low-temperature breakdown.
@@ Appendix B Feynman–Dyson nonconvergence
We reportedHirata et al. (2024) pervasive divergences of a Feynman–Dyson perturbation expansion of the self-energy in many frequency domains. Therefore, apart from the Kohn–Luttinger-type nonconvergence of the thermal self-energy caused by anomalous-diagram terms, the zero-temperature self-energy will already be divergent or at least excessively erroneous for low- and high-lying states. This is alleviated (or hidden from view) by the frequency-independent approximation inherent in the ansatz of thermal quasi-particle theory, but in a solid, whose self-energy forms continuous energy bands, we can no longer hope to be able to avoid this deep-rooted pathology.Hirata et al. (2024)
In a solid-state implementation of thermal QP(2) theory (which is underway in our laboratory), therefore, we will have to introduce active orbitals (or active energy bands) and evaluate only some portions of $\langle E^{(2)}\rangle$ and $\Sigma_{pp}^{(2)}$ involving them. These active orbitals are the ones whose zeroth-order energies fall within the so-called “central overlapping bracket” where the Feynman–Dyson perturbation expansion is guaranteed to have a nonzero radius of convergence.Hirata et al. (2024) Comparative performance of this method for solids will be presented in the future.